A rectangular lawn has dimensions 30 metres by 40 metres. A strip of uniform width is to be mowed around the perimeter.How wide should the strip be so that exactly half of the lawn gets mowed ?
Answer by checkley71(8403)

30*40=2(30-2X)(40-2X)
1200=2(1200-80X-60X+4X^2)
1200=2(1200-140X+4X^2)
1200=2400-280X+8X^2
8X^2-280X+2400-1200=0
8X^2-280X+1200
8(X^2-35X+150)
8(X-5)(X-30)=0
X-5=0
X=5 ANSWER FOR THE WIDTH OF THE STRIP MOWED TO REDUCE THE LAWN TO 50%.
PROOF
(30-2*5)(40-2*5)=1200/2
20*30=600
600=600